OpenStudy (anonymous):

A portion of fossilized bone has lost 97% of its original amount of carbon-14, which has a half-life of 5730 years. Assuming exponential decay, estimate the age of the fossil. Can someone explain how to solve this?

OpenStudy (anonymous):

(3/100)P=P(1/2)^(t/5730) I don't understand where the 3/100 came from..

OpenStudy (anonymous):

hmm that would make sense....but if I didn't know that I would of put 97/100..

OpenStudy (anonymous):

dN/dt = - k N dN/N = - k dt ln(N) = -k t N(t) = N(0) exp(-k t) (N(5730)/ N(0) = 0.5 = exp(- k 5730) ln(0.5)= - 5730 k k = 0.693/5730 97% lost means 3% left 0.03 = exp(- t 0.693/5730) t = -(5730/0.693) ln (0.03) t = 28994 years

OpenStudy (anonymous):

oh nvm I get it thanksXD

OpenStudy (anonymous):

Very good.